OpenStudy (anonymous):

@aajugdar The base of a particular parallelogram is 8 times its height. If the base and height are doubled, the new area is 384 square inches greater than the area of the original parallelogram. Find the base and height of the original parallelogram. A. Write the equation for the problem. B. Solve. Show all work. C. Name the length of the base and height of the original parallelogram. Remember the labels.

OpenStudy (anonymous):

Area = b*h let h be height of original, that makes the base 8h (8 times the base) Let h * 8h=A for area of original parallelogram 8h^2=A The 2nd parallelogram the height is 2h and the base is 16h (doubled) and the area is 384 more than the original area so 2h*16h=A+384 32h^2 = A+384 32h^2-384=A Substitute 8h^2 from the first formula for A in the second formula 32h^2-384=8h^2 24h^2=384 h^2=384/24 h^2=16 h=4 original height is 4 which makes the base 32

OpenStudy (anonymous):

thats correct

OpenStudy (anonymous):

Thnks